# HQ-VAE: Hierarchical Discrete Representation Learning with Variational Bayes

### v1

@@ 1 Introduction
Vector quantization is usually achieved with a vector quantized variational autoencoder (VQ-VAE) (van den Oord et al., 2017). In the original VQ-VAE, an input is first encoded and quantized with code vectors, which produces a discrete representation of the encoded feature. The discrete representation is then decoded to the data space to recover the original input. Subsequent developments incorporated a hierarchical structure into the discrete latent space to achieve high-fidelity reconstructions. In particular, Razavi et al. (2019) developed VQ-VAE into a hierarchical model, called VQ-VAE-2. In this model, multi-resolution discrete latent representations are used to extract local and global information from the target data. Another type of hierarchical discrete representation, called residual quantization (RQ), was proposed to reduce the gap between the feature maps before and after the quantization process (Zeghidour et al., 2021; Lee et al., 2022a).
@@ 2 Background
First, we revisit VQ-VAE and its extensions to hierarchical latent models. Then, we review SQ-VAE, which serves as the foundation framework of HQ-VAE.
2.1 VQ-VAE
To discretely represent observations $\bm{x}\in\mathbb{R}^{D}$, a codebook $\bm{B}$ is used that consists of finite trainable code vectors $\{\bm{b}_{k}\}_{k=1}^{K}$ ($\bm{b}_{k}\in\mathbb{R}^{d_{b}}$). A discrete latent variable $\bm{Z}$ is constructed to be in a $d_{z}$-tuple of $\bm{B}$, i.e., $\bm{Z}\in\bm{B}^{d_{z}}$, which is later decoded to generate data samples. A deterministic encoder and decoder pair is used to connect the observation and latent representation, where the encoder maps $\bm{x}$ to $\bm{Z}$ and the decoder recovers $\bm{x}$ from $\bm{Z}$ by using a decoding function $\bm{f}_{\bm{\theta}}:\mathbb{R}^{d_{b}\times d_{z}}\to\mathbb{R}^{D}$. An encoding function, denoted as $\bm{G}_{\bm{\phi}}:\mathbb{R}^{D}\to\mathbb{R}^{d_{b}\times d_{z}}$, and a deterministic quantization operator are used together as the encoder. The encoding function first maps $\bm{x}$ to $\hat{\bm{Z}}\in\mathbb{R}^{d_{b}\times d_{z}}$; then, the quantization operator finds the nearest neighbor of $\hat{\bm{z}}_{i}$ for $i\in[d_{z}]$, i.e., $\bm{z}_{i}=\operatorname*{arg\,min}_{\bm{b}_{k}}\|\hat{\bm{z}}_{i}-\bm{b}_{k}\|_{2}^{2}$. The trainable components (encoder, decoder, and codebook) are learned by minimizing the objective,
@@ 2.1 VQ-VAE
VQ-VAE-2. To model local and global information separately, VQ-VAE-2 uses a hierarchical structure for vector quantization (Razavi et al., 2019). The model consists of multiple levels of latents so that the top levels have global information, while the bottom levels focus on local information, conditioned on the top levels. The training of the model follows the same scheme as the original VQ-VAE (e.g., stop-gradient, EMA update, and deterministic quantization).
@@
RQ-VAE. RQ provides a finer approximation of $\bm{Z}$ by taking into account information on quantization gaps (residuals) (Zeghidour et al., 2021; Lee et al., 2022a). With RQ, $L$ code vectors are assigned to each vector $\bm{z}_{i}$ ($i\in[d_{z}]$), instead of increasing the codebook size $K$. To make multiple assignments, RQ repeatedly quantizes the target feature and computes quantization residuals, denoted as $\bm{R}_{l}$. Namely, the following procedure is repeated $L$ times, starting with $\bm{R}_{0}=\hat{\bm{Z}}$: $\bm{z}_{l,i}=\operatorname*{arg\,min}_{\bm{b}_{k}}\|\bm{r}_{l-1,i}-\bm{b}_{k}\|_{2}^{2}\quad\text{and}\quad\bm{R}_{l}=\bm{R}_{l-1}-\bm{Z}_{l}$. By repeating RQ, the discrete representation can be refined in a coarse-to-fine manner. Finally, RQ discretely approximates the encoded variable as $\hat{\bm{Z}}\approx\sum_{l=1}^{L}\bm{Z}_{l}$, where the conventional VQ is regarded as a special case of RQ with $L=1$.
2.2 SQ-VAE
@@
SQ-VAE (Takida et al., 2022b) also has deterministic encoding/decoding functions and a trainable codebook like the above autoencoders. However, unlike the deterministic quantization scheme of VQ and RQ, SQ-VAE follows the variational Bayes framework and has an SQ procedure for the encoded features. More precisely, it defines a stochastic dequantization process $p_{s^{2}}(\tilde{\bm{z}}_{i}|\bm{Z})=\mathcal{N}(\tilde{\bm{z}}_{i};\bm{z}_{i},s^{2}\bm{I})$, which converts a discrete variable $\bm{z}_{i}$ into a continuous one $\tilde{\bm{z}}_{i}$ by adding Gaussian noise with a learnable variance $s^{2}$. By Bayes’ rule, this process is associated with the reverse operation, i.e., SQ, which is given by $\hat{P}_{s^{2}}(\bm{z}_{i}=\bm{b}_{k}|\tilde{\bm{Z}})\propto\exp\left(-\frac{\|\tilde{\bm{z}}_{i}-\bm{b}_{k}\|_{2}^{2}}{2s^{2}}\right)$. Thanks to this variational framework, the degree of stochasticity in the quantization scheme becomes adaptive. This allows SQ-VAE to benefit from the effect of self-annealing, where the SQ process gradually approaches a deterministic one as $s^{2}$ decreases. This generally improves the efficiency of codebook usage.
SQ-VAE vs. dVAE. Sønderby et al. (2017) proposed a discrete latent VAE with a codebook lookup. Ramesh et al. (2021) took the same approach to train a VAE equipped with a codebook for text-to-image generation; they called their model discrete VAE (dVAE). In dVAE, a stochastic posterior categorical distribution under which code vectors are assigned is directly modeled by the encoder as $Q(\bm{z}_{i}=\bm{b}_{k}|\bm{x})\propto\exp(g_{\bm{\phi},k}(\bm{x}))$, where $\bm{g}_{\bm{\phi}}:\mathbb{R}^{D}\to\mathbb{R}^{K}$. This modeling enables the original sample $\bm{x}$ to be encoded into a set of codes $\bm{Z}$ in a probabilistic way. However, such an index-domain modeling cannot incorporate the codebook geometry explicitly into the posterior modeling. In contrast, VQ/SQ-based methods allow us to model the posterior distribution with vector operations (van den Oord et al., 2017; Lee et al., 2022a; You et al., 2022). Another benefit of VQ/SQ-VAE is that they can evaluate the reconstruction errors coming from the discretization with the quantization errors (Dhariwal et al., 2020). Note that, in what follows, we restrict the scope to latent generative models involving vector quantization operators.
(a) Overview of HQ-VAE.
(b) First top-down layer.
(c) Injected top-down layer.
@@
Figure 1: (a) HQ-VAE consists of bottom-up and top-down paths. Red arrows are for the approximated posterior. The objective function takes the Kullback–Leibler divergence of posterior and prior (in the blue box). (b) First layer for top-down path. (c)-(d) There are two types of top-down layer: injected top-down and residual top-down. An HQ-VAE that consists entirely of the injected (residual) top-down layers is analogous to VQ-VAE-2 (RQ-VAE).
@@ 3 Hierarchically quantized VAE
To achieve a hierarchical discrete representation of depth $L$, we first divide up the discrete latent variables into $L$ groups of discrete latent variables, which are denoted as $\bm{Z}_{1:L}:=\{\bm{Z}_{l}\}_{l=1}^{L}$. For each $l\in[L]$, we further introduce a trainable codebook, $\bm{B}_{l}:=\{\bm{b}_{k}^{l}\}_{k=1}^{K_{l}}$, consisting of $K_{l}$ $d_{b}$-dimensional code vectors, i.e., $\bm{b}_{k}^{l}\in\mathbb{R}^{d_{b}}$ for $k\in[K_{l}]$. The variable $\bm{Z}_{l}$ is represented as a $d_{l}$-tuple of the code vectors in $\bm{B}_{l}$; namely, $\bm{Z}_{l}\in\bm{B}_{l}^{d_{l}}$. Similarly to conventional VAEs, the latent variable of each group is assumed to follow a pre-defined prior mass function. We set the prior as an i.i.d. uniform distribution, defined as $P(\bm{z}_{l,i}=\bm{b}_{k})=1/K_{l}$ for $i\in[d_{l}]$. The probabilistic decoder is set to a normal distribution with a trainable isotropic covariance matrix, $p_{\bm{\theta}}(\bm{x}|\bm{Z}_{1:L})=\mathcal{N}(\bm{x};\bm{f}_{\bm{\theta}}(\bm{Z}_{1:L}),\sigma^{2}\bm{I})$ with a decoding function $\bm{f}_{\bm{\theta}}:\mathbb{R}^{d_{b}\times d_{1}}\oplus\cdots\oplus\mathbb{R}^{d_{b}\times d_{L}}\to\mathbb{R}^{D}$. To generate instances, it decodes latent variables sampled from the prior. Here, the exact evaluation of $P_{\bm{\theta}}(\bm{Z}_{1:L}|\bm{x})$ is required to train the generative model with the maximum likelihood. However, this is intractable in practice. Thus, we introduce an approximated posterior on $\bm{Z}_{1:L}$ given $\bm{x}$ and derive the evidence lower bound (ELBO) for maximization instead.
Inspired by the work on hierarchical Gaussian VAEs (Sønderby et al., 2016; Vahdat & Kautz, 2020; Child, 2021), we give HQ-VAE bottom-up and top-down paths, as shown in Figure 1. The approximated posterior has a top-down structure ($\bm{Z}_{1}\to\bm{Z}_{2}\to\cdots\to\bm{Z}_{L}$). The bottom-up path first generates features from $\bm{x}$ as $\bm{H}_{\bm{\phi}}^{r}(\bm{x})$ at different resolutions ($r\in[R]$). The latent variable in each group on the top-down path is processed from $\bm{Z}_{1}$ to $\bm{Z}_{L}$ in that order by taking $\bm{H}_{\bm{\phi}}^{r}(\bm{x})$ into account. To do so, two features, one extracted by the bottom-up path ($\bm{H}_{\bm{\phi}}^{r}(\bm{x})$) and one processed on the higher layers of the top-down path ($\bm{Z}_{1:l-1}$), can be fed to each layer and used to estimate the $\bm{Z}_{l}$ corresponding to $\bm{x}$, which we define as $\hat{\bm{Z}}_{l}=\bm{G}_{\bm{\phi}}^{l}(\bm{H}_{\bm{\phi}}^{r}(\bm{x}),\bm{Z}_{1:l-1})$. The $l$th group $\bm{Z}_{l}$ has a unique resolution index $r$; we denote it as $r(l)$. For simplicity, we will ignore $\bm{H}_{\bm{\phi}}^{r}$ in $\hat{\bm{Z}}_{l}$ and write $\hat{\bm{Z}}_{l}=\bm{G}_{\bm{\phi}}^{l}(\bm{x},\bm{Z}_{1:l-1})$. The design of the encoding function $\bm{G}_{\bm{\phi}}^{l}$ lead us to a different modeling of the approximated posterior. A detailed discussion of it is presented in the next section.
@@ 4 Instances of HQ-VAE
Now that we have established the overall framework of HQ-VAE, we will consider two special cases with different top-down layers: injected top-down or residual top-down. We will derive two instances of HQ-VAE that consist only of the injected top-down layer or the residual top-down layer, which we will call SQ-VAE-2 and RSQ-VAE by analogy to VQ-VAE-2 and RQ-VAE. These two layers can be combinatorially used to define a hybrid model of SQ-VAE-2 and RSQ-VAE, as explained in Appendix C. Note that the prior distribution (Equation (4)) is identical across all instances.
@@ 4.2.1 SQ-VAE-2
where $H(\cdot)$ indicates the entropy of the probability mass function and constant terms have been omitted. The derivation of Equation (6) is in Appendix B. The objective function (6) consists of a reconstruction term and regularization terms for $\bm{Z}_{1:L}$ and $\tilde{\bm{Z}}_{1:L}$. The expectation w.r.t. the probability mass function $\hat{P}_{s_{l}^{2}}(\bm{z}_{l,i}=\bm{b}_{k}|\tilde{\bm{Z}}_{l})$ can be approximated with the corresponding Gumbel-softmax distribution (Maddison et al., 2017; Jang et al., 2017) in a reparameterizable manner.

### v2

@@ 1 Introduction
Vector quantization is usually achieved with a vector quantized variational autoencoder (VQ-VAE) (van den Oord et al., 2017). Based on the construction of van den Oord et al. (2017), inputs are first encoded and quantized with code vectors, which produces a discrete representation of the encoded feature. The discrete representation is then decoded to the data space to recover the original input. Subsequent developments incorporated a hierarchical structure into the discrete latent space to achieve high-fidelity reconstructions. In particular, Razavi et al. (2019) developed VQ-VAE into a hierarchical model, called VQ-VAE-2. In this model, multi-resolution discrete latent representations are used to extract local (e.g., texture in images) and global (e.g., shape and geometry of objects in images) information from the target data. Another type of hierarchical discrete representation, called residual quantization (RQ), was proposed to reduce the gap between the feature maps before and after the quantization process (Zeghidour et al., 2021; Lee et al., 2022a).
@@ 2 Background
We first review VQ-VAE and its extensions to hierarchical latent models. Then, we revisit SQ-VAE, which serves as the foundation framework of HQ-VAE.
VQ-VAE. To discretely represent observations $\bm{x}\in\mathbb{R}^{D}$, a codebook $\bm{B}$ is used that consists of finite trainable code vectors $\{\bm{b}_{k}\}_{k=1}^{K}$ ($\bm{b}_{k}\in\mathbb{R}^{d_{b}}$). A discrete latent variable $\bm{Z}$ is constructed to be in a $d_{z}$-tuple of $\bm{B}$, i.e., $\bm{Z}\in\bm{B}^{d_{z}}$, which is later decoded to generate data samples. A deterministic encoder and decoder pair is used to connect the observation and latent representation, where the encoder maps $\bm{x}$ to $\bm{Z}$ and the decoder recovers $\bm{x}$ from $\bm{Z}$ by using a decoding function $\bm{f}_{\bm{\theta}}:\mathbb{R}^{d_{b}\times d_{z}}\to\mathbb{R}^{D}$. An encoding function, denoted as $\bm{G}_{\bm{\phi}}:\mathbb{R}^{D}\to\mathbb{R}^{d_{b}\times d_{z}}$, and a deterministic quantization operator are used together as the encoder. The encoding function first maps $\bm{x}$ to $\hat{\bm{Z}}\in\mathbb{R}^{d_{b}\times d_{z}}$; then, the quantization operator finds the nearest neighbor of $\hat{\bm{z}}_{i}$ for $i\in[d_{z}]$, i.e., $\bm{z}_{i}=\operatorname*{arg\,min}_{\bm{b}_{k}}\|\hat{\bm{z}}_{i}-\bm{b}_{k}\|_{2}^{2}$. The trainable components (encoder, decoder, and codebook) are learned by minimizing the objective,
@@
VQ-VAE-2. Razavi et al. (2019) incorporated hierarchical structure into the discrete latent space in VQ-VAE to model local and global information separately. The model consists of multiple levels of latents so that the top levels have global information, while the bottom levels focus on local information, conditioned on the top levels. The training of the model follows the same scheme as the original VQ-VAE (e.g., stop-gradient, EMA update, and deterministic quantization).
@@
RQ-VAE. As an extension of VQ, RQ was proposed to provide a finer approximation of $\bm{Z}$ by taking into account information on quantization gaps (residuals) (Zeghidour et al., 2021; Lee et al., 2022a). With RQ, $L$ code vectors are assigned to each vector $\bm{z}_{i}$ ($i\in[d_{z}]$), instead of increasing the codebook size $K$. To make multiple assignments, RQ repeatedly quantizes the target feature and computes quantization residuals, denoted as $\bm{R}_{l}$. Namely, the following procedure is repeated $L$ times, starting with $\bm{R}_{0}=\hat{\bm{Z}}$: $\bm{z}_{l,i}=\operatorname*{arg\,min}_{\bm{b}_{k}}\|\bm{r}_{l-1,i}-\bm{b}_{k}\|_{2}^{2}\quad\text{and}\quad\bm{R}_{l}=\bm{R}_{l-1}-\bm{Z}_{l}$. By repeating RQ, the discrete representation can be refined in a coarse-to-fine manner. Finally, RQ discretely approximates the encoded variable as $\hat{\bm{Z}}\approx\sum_{l=1}^{L}\bm{Z}_{l}$, where the conventional VQ is regarded as a special case of RQ with $L=1$.
@@
SQ-VAE. Takida et al. (2022b) proposed a variational Bayes framework for learning the VQ-VAE components to mitigate the issue of codebook collapse. The resulting model, SQ-VAE, also has deterministic encoding/decoding functions and a trainable codebook like the above autoencoders. However, unlike the deterministic quantization schemes of VQ and RQ, SQ-VAE adopts stochastic quantization (SQ) for the encoded features to approximate the categorical posterior distribution $P(\bm{Z}|\bm{x})$. More precisely, it defines a stochastic dequantization process $p_{s^{2}}(\tilde{\bm{z}}_{i}|\bm{Z})=\mathcal{N}(\tilde{\bm{z}}_{i};\bm{z}_{i},s^{2}\bm{I})$, which converts a discrete variable $\bm{z}_{i}$ into a continuous one $\tilde{\bm{z}}_{i}$ by adding Gaussian noise with a learnable variance $s^{2}$. By Bayes’ rule, this process is associated with the reverse operation, i.e., SQ, which is given by $\hat{P}_{s^{2}}(\bm{z}_{i}=\bm{b}_{k}|\tilde{\bm{Z}})\propto\exp\left(-\frac{\|\tilde{\bm{z}}_{i}-\bm{b}_{k}\|_{2}^{2}}{2s^{2}}\right)$. Thanks to this variational framework, the degree of stochasticity in the quantization scheme becomes adaptive. This allows SQ-VAE to benefit from the effect of self-annealing, where the SQ process gradually approaches a deterministic one as $s^{2}$ decreases. This generally improves the efficiency of codebook usage. Other related works in the literature of discrete posterior modeling are found in Appendix A.
@@
Figure 1: (a) HQ-VAE consists of bottom-up and top-down paths. Red arrows represent the approximated posterior (Equation (3)). This process consists of the following three steps: the bottom-up path generates the features $\{\bm{H}_{\bm{\phi}}^{r}(\bm{x})\}_{r=1}^{R}$ at different resolutions; the top-down path fuses each of them with the information processed in the top-down path into $\tilde{\bm{Z}}_{1:L}$; the auxiliary variable at the $l$th layer, $\tilde{\bm{Z}}_{l}$, is quantized into $\bm{Z}_{l}$ based on the codebook $\bm{B}_{l}$. Lastly, to complete the autoencoding process, all the quantized latent variables in the top-down path, $\bm{Z}_{1:L}$, are fed to the decoder $\bm{f}_{\bm{\theta}}$, recovering the original data sample. The objective function takes the Kullback–Leibler divergence of posterior and prior (in the blue box). (b) First layer for top-down path. Any HQ-VAE should include this layer as its top layer. An HQ-VAE that includes only this layer in its top-down path reduces to an SQ-VAE. (c)-(d) Two types of top-down layers are proposed: injected top-down and residual top-down layers. Injected top-down layer newly injects the higher-resolution features generated by the bottom-up path into the latent variables in the previous top-down layers, resulting in a higher resolution (i.e., larger dimension) of $\tilde{\bm{Z}}_{l}$ and $\bm{Z}_{l}$ compared to $\bm{Z}_{1:l-1}$. Residual top-down layer does not incorporate any new features from the bottom-up path but aims to refine the processed latent variables in top-down path with additional code vectors for representing the quantization residual from the preceding top-down layer. As a result, it maintains the same dimensionality for $\bm{Z}_{l}$ as in the previous layer (i.e., $\bm{Z}_{l-1}$). An HQ-VAE that consists entirely of the injected (residual) top-down layers is analogous to VQ-VAE-2 (RQ-VAE), which is presented in Section 4.2.1 (4.3.1).
@@ 3 Hierarchically quantized VAE
To achieve a hierarchical discrete representation of depth $L$, we first introduce $L$ groups of discrete latent variables, which are denoted as $\bm{Z}_{1:L}:=\{\bm{Z}_{l}\}_{l=1}^{L}$. For each $l\in[L]$, we further introduce a trainable codebook, $\bm{B}_{l}:=\{\bm{b}_{k}^{l}\}_{k=1}^{K_{l}}$, consisting of $K_{l}$ $d_{b}$-dimensional code vectors, i.e., $\bm{b}_{k}^{l}\in\mathbb{R}^{d_{b}}$ for $k\in[K_{l}]$. The variable $\bm{Z}_{l}$ is represented as a $d_{l}$-tuple of the code vectors in $\bm{B}_{l}$; namely, $\bm{Z}_{l}\in\bm{B}_{l}^{d_{l}}$. Similarly to conventional VAEs, the latent variable of each group is assumed to follow a pre-defined prior mass function. We set the prior as an i.i.d. uniform distribution, defined as $P(\bm{z}_{l,i}=\bm{b}_{k})=1/K_{l}$ for $i\in[d_{l}]$. The probabilistic decoder is set to a normal distribution with a trainable isotropic covariance matrix, $p_{\bm{\theta}}(\bm{x}|\bm{Z}_{1:L})=\mathcal{N}(\bm{x};\bm{f}_{\bm{\theta}}(\bm{Z}_{1:L}),\sigma^{2}\bm{I})$ with a decoding function $\bm{f}_{\bm{\theta}}:\mathbb{R}^{d_{b}\times d_{1}}\oplus\cdots\oplus\mathbb{R}^{d_{b}\times d_{L}}\to\mathbb{R}^{D}$. To generate instances, it decodes latent variables sampled from the prior. Here, the exact evaluation of $P_{\bm{\theta}}(\bm{Z}_{1:L}|\bm{x})$ is required to train the generative model with the maximum likelihood. However, this is intractable in practice. Thus, we introduce an approximated posterior on $\bm{Z}_{1:L}$ given $\bm{x}$ and derive the evidence lower bound (ELBO) for maximization instead.
Inspired by the work on hierarchical Gaussian VAEs (Sønderby et al., 2016; Vahdat & Kautz, 2020; Child, 2021), we introduce HQ-VAE bottom-up and top-down paths, as shown in Figure 1. The approximated posterior has a top-down structure ($\bm{Z}_{1}\to\bm{Z}_{2}\to\cdots\to\bm{Z}_{L}$). The bottom-up path first generates features from $\bm{x}$ as $\bm{H}_{\bm{\phi}}^{r}(\bm{x})$ at different resolutions ($r\in[R]$). The latent variable in each group on the top-down path is processed from $\bm{Z}_{1}$ to $\bm{Z}_{L}$ in that order by taking $\bm{H}_{\bm{\phi}}^{r}(\bm{x})$ into account. To do so, two features, one extracted by the bottom-up path ($\bm{H}_{\bm{\phi}}^{r}(\bm{x})$) and one processed on the higher layers of the top-down path ($\bm{Z}_{1:l-1}$), can be fed to each layer and used to estimate the $\bm{Z}_{l}$ corresponding to $\bm{x}$, which we define as $\hat{\bm{Z}}_{l}=\bm{G}_{\bm{\phi}}^{l}(\bm{H}_{\bm{\phi}}^{r}(\bm{x}),\bm{Z}_{1:l-1})$. The $l$th group $\bm{Z}_{l}$ has a unique resolution index $r$; we denote it as $r(l)$. For simplicity, we will ignore $\bm{H}_{\bm{\phi}}^{r}$ in $\hat{\bm{Z}}_{l}$ and write $\hat{\bm{Z}}_{l}=\bm{G}_{\bm{\phi}}^{l}(\bm{x},\bm{Z}_{1:l-1})$. The design of the encoding function $\bm{G}_{\bm{\phi}}^{l}$ lead us to a different modeling of the approximated posterior. A detailed discussion of it is presented in the next section.
@@ 4 Instances of HQ-VAE
Now that we have established the overall framework of HQ-VAE, we will consider two special cases with different top-down layers: injected top-down or residual top-down. We will derive two instances of HQ-VAE that consist only of the injected top-down layer or the residual top-down layer, which we will call SQ-VAE-2 and RSQ-VAE by analogy to VQ-VAE-2 and RQ-VAE. Specifically, SQ-VAE-2 and RSQ-VAE share the same architectural structure as VQ-VAE-2 and RQ-VAE, respectively, but the differences lie in their training strategies, as discussed in Sections 4.2.2 and 4.3.2. Furthermore, these two layers can be combinatorially used to define a hybrid model of SQ-VAE-2 and RSQ-VAE, as explained in Appendix D. Note that the prior distribution (Equation (4)) is identical across all instances.
@@ 4.2.1 SQ-VAE-2
where $H(\cdot)$ indicates the entropy of the probability mass function and constant terms have been omitted. The derivation of Equation (6) is in Appendix C. The objective function (6) consists of a reconstruction term and regularization terms for $\bm{Z}_{1:L}$ and $\tilde{\bm{Z}}_{1:L}$. The expectation w.r.t. the probability mass function $\hat{P}_{s_{l}^{2}}(\bm{z}_{l,i}=\bm{b}_{k}|\tilde{\bm{Z}}_{l})$ can be approximated with the corresponding Gumbel-softmax distribution (Maddison et al., 2017; Jang et al., 2017) in a reparameterizable manner.